These are the news starting 01 Jan 2009
Click here for the news of the previous years.
|31 Dec 09||Made overplot of Uccle and meteoLCD total ozone column graphs.(see DOBSON page)|
|24 Dec 09||Recalibration and maintenance of CO2 and O3 sensors.|
|21 Dec 09||
1. repaired ultrasonic anemometer back into action (13:45).
There may be some problems during the next day.
2. Calibration and maintenance of the NO/NO2 sensor
3. The cup anemometer was out of action during the last cold days: it was stuck frozen!
|06 Dec 09||Uploaded nov09.dat to data archive|
|05 Dec 09||Maintenance and recalibration of CO2 and O3 sensors.|
|02 Dec 09||I am back from 3 weeks spent in the USA. Regular contributions will resume.|
|10 Nov 09||Maintenance and calibration work.|
|04 Nov 09||Uploaded oct09.dat to data archive|
|02 Nov 09||Start of the Klima2009 on-line conference organized by the Faculty of Applied Sciences, University of Hamburg. Ernst-Georg Beck and myself have submitted a peer-reviewed paper on CO2 background estimation that attracts a lot of visitors.|
|28 Oct 09||Ultrasonic anemometer has been dismounted to be sent back for repair. Look here for two pictures (PIC1, PIC2); the forst showing me doing some acrobatics, the second showing technician Jean-Paul Klein who helped.|
|17 Oct 09||
1. Maintenance and recalibration of gaz sensors.
2. Webcam proxy computer was dead (probably since a good many days). Checked and restarted: OK.
|05 Oct 09||Uploaded sep09.dat to data archive.|
|02 Oct 09||New compressor mounted in LNI zero-air generator. CO2 and NOx sensors fitted back to zero-air generator.|
|23 Sep 09||
Maintenance and recalibration work by Envitec. The LNI air zero
generator is out of order (compressor breakdown); we switched
back to classical scrubbers.
The whole teflon tubing and manifolds of the air admission system has been thoroughly cleaned...there was much black grime to remove!
|15 Sep 09||Microtops #5375 is back from calibration at Mauna Loa and Solar Light Co. The new Dobson readings will be given in bold purple. I will make measurements for some time with the 2 instruments #3012 and #5375 (who do not use exactly the same wavelengths!), and a comparison report written asap.|
|09 Sep 09||Uploaded aug01.dat to data archive.|
|20 Aug 09||Maintenance work and recalibration of the gas sensors.|
|12 Aug 09||Lightning computer working again! Cause of malfunction remains mysterious.|
|01Aug 09||Uploaded aug09.dat file into data archive.|
|31 Jul 09||The server meteo01.lcd.lu who hosts the lightning images is down. Sorry about this! Repair not possible before 17th August.|
|22 Jul 09||Maintenance work and recalibration of the gas sensors.|
|12 Jul 09||
1. Uploaded jun09.dat to data archive
2. During the past weeks, the ultrasonic anemometer (sensor AN3) had intermittent failures (extremely high spikes). Use data from cup anemometer AN1 in case of doubt. The relationship between both sensors is:
AN3 = 0.923*AN1 +0.345 (R2=0.97) or the more simple linear relationship:
AN3 = 1.042*AN1 (R2=0.95)
These linear (affine) fits have been calculated on more than 2200 readings from 22May09 to 12Jul09
Since over a week, the ultrasonic anemometer is ok again...
|02 Jul 09||There is a serious problem with the ultra-sonic wind sensor (ch.4, air_velo). Please use the cup-anemometer readings fron ch.61, AirSpeed.) I will adapt the graphs asap.|
|29 Jun 09||
1. Recalibrated CO2 sensor.
2. Edited the flashtype pages: flash-icons have slighlty be modified to increase visiblity (new, new, old and older categories)
|23 Jun 09||
The Microtops instrument #5375 is back to Solar Light for
recalibration during Aug.09. This will be done on the Mauna Loa
site in Hawaï. Readings from older #3012 instrument are
adjusted to the previous by a multiplier of 1.02955 (found from
a comparison study using both instruments done in May/June):
It should be noted that the recalibration is paid by sponsor Dr. Nic Kahn (a former student of the LCD and also a contributor to our Computarium). Many thanks to Dr. Kahn for allowing this expensive but necessary recalibration!
|22 Jun 09||I submitted together with Ernst-Georg Beck (author of a famous book:on historic CO2 measurements, see article here) a paper on our CO2/windspeed study to Klima2009, an organisation of the Hamburg University of Applied Sciences (ours is one of the 150 papers accepted). Klima2009 is a series of on-line research conferences (2nd to the 6th November 2009), presenting peer-reviewed papers. Our submission has been accepted; let us see what will be the fate of the full blown paper (sorry: not publicly available before the conference!)|
|13 Jun 09||
1. Corrected a faulty link on the main page of http://meteo.lcd.lu
to access the webcam(port 8080 was missing).
2. Recalibration of CO2 sensor
3. meteoLCD is given as the premier (and single) source of UVI data for Luxembourg by the "Fondation Luxembourgeoise contre le Cancer" in their Jun09 brochure "Safer Sun" (see page 4)
|12 Jun 09||Maintenance work by Envitec|
|05 Jun 09||Uploaded may09.dat to data archive.|
|22 May 09||There was a severe multi-hour power outage for all the town of Diekirch during the night. The main logger stopped logging, because it's internal batteries had run flat.. Repairs and data downloaded now (07:45 UTC) .Logging will restart at 08:00 UTC..|
|21 May 09||I updated my paper on the trend on atmospheric optical thickness by including findings from a new paper of R. Vautard at al. (Nature Geoscience): this new paper confirms the meteoLCD observation of a reversal of the year-long downtrend of the atmospheric turpitude into an increase, starting around 2000-2003.|
|20 May 09||
Finished maintenance and calibration work with AC31M (NOx)
The new Lightning2000 software seems to be immune from memory creep.
|11 May 09||There was a problem with the lightning computer (ftp buffer filling up). I installed today a new version of the Lightning2000 software, that, according to Aninoquisi should have much better nearby flash detection.|
|07 May 09||
1. Maintenance and calibration work on O3 and CO2 sensors.
2. Lightning computer was terminally ill. Replaced.
|06 May 09||Replaced temperature/humidity sensing module (Humicor S6000). Humidity readings were unreliable since some time, and sensor did not recover good health.|
|03 May 09||Uploaded apr09 data into archive.|
|03 Apr 09||Maintenance work by Envitec.|
|03 Mar 09||
1. Repaired a silly mistake that blocked access to webcam
2. Uploaded Feb09 data into archive.
|21 Feb 09||I had a serious computer problem when replacing the webcam proxy computer with a new one. Due to an obscure (?) rights miss match the proxy didn't work properly (during the last days) when called from outside of our domain. This is fixed now.|
|14 Feb 09||Maintenance work and recalibration on all gas sensors.. Large correction of CO2 sensor: zero was possibly -22 ppm too low. Be careful with readings from about 01Feb. on, they might be too low. (last calibration was done by Envitec on 22Jan09). I redid zero and span calibration two times, values are stable.|
|12 Feb 09||
I made some heavy changes on the computer hardware. Three older
PC's have been replaced by very small fan-less SUMICOM
computers (which were on use for one year at the Infoscreens
stations of the LCD). The following computers have been
replaced: logger computer (was an old NEC), radioactivity
computer (a clone) and the Davis backup weatherstation computer
(a real antique Siemens (not FSC) machine). All these new little
gems run W2K SP4. I added the "speedfan" utility and
"rockettime" to guarantee a correct clock.
The Davis machine was replaced today; I hope I got all these configurations right.
The Linux box running the proxy for the webcam will be replaced asap. I will probably switch to openSUSE 11.1.
|08 Feb 09||
1. Uploaded jan09.dat to data archive.
2. Lightning data not available for the moment because of ongoing work on the server (you won't miss much!)
|01 Feb 09||I wrote a short paper on CO2 (and other gases) pattern during the 4 days long cold spell from 09 to 12th Jan. 2009. See "Papers and Comments" section or here.|
|23 Jan 09||
1. The antique logger computer has been replaced by a younger
Sumicom fan-less pico computer, running well under W2K SP4.
2. Idem for the radioactivity computer (some hikups with WinXP, so I changed to W2K SP4, which is ok.
|22 Jan 09||Maintenance work by Envitec.|
|21 Jan 09||Black Wednesday today! Two computers went bust: the main logger computer and the radioactivity computer (both were very antique!). I replaced them with fanless pico-computers running W2K SP4 (for the logger computer: he must execute a very old Turbopascal compiled program) and XP Pro (radioactivity computer running the RS03control software from Claude Baumann)|
|17 Jan 09||
Update of the local trends 1998 to 2008. DTR
trend will be done asap (needs more calculations).
Strong negative trends for all solar related parameters!
|01 Jan 09||
1. The usual new year problem with the 7 days
graphs.. as the logger file does not contain the year, Gnuplot
gets confused the 1rst week of every new year. This will self
correct the 7th January.
2. Uploaded dec08.dat and 2008.zip files to the data archive